OpenStudy (anonymous):

Write an equation in standard form with integer coefficents for the line with slope 14/19going through the point (-1,-6)

OpenStudy (anonymous):

slope 14/19 going through the point (-1,-6) -> y + 6 = (14/19) ( x + 1) -> 19( y + 6) = 14 ( x + 1) => 14x - 19 y = 100

OpenStudy (anonymous):

Sorry about that. I didn't put the negative in in the last step. :-( -6 = (14/19)*-1 +b b = -100/19 y=mx+b m=14/19 b=-100/19 y=(14/19)*x -100/19